ZEKE SPECTROSCOPY OF Al(C_{5}H_{5}N): A STUDY OF METAL-σ INTERACTION WITH A HETERO-CYCLIC AROMATIC LIGAND

\underline{XU WANG}, SHENGGANG LI, JASON F. FULLER, PARAGRANJITA BHOWMIK, BRADFORD R. SOHNLEIN, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

A few years ago, a computational study was reported on cation-π interactions in Al^{+} complexes with a series of aromatic compounds^{a}. However, for Al^{+}(C_{5}H_{5}N), the calculations indicated that Al^{+} was σ-bound to the nitrogen in the ground electronic state. If Al^{+} was π-bound to the ring, the potential energy surface was very flat and no minimum was located. Our photoelectron measurements on Al/Al^{+}-pyridine vibrations agree with the theoretical predictions. The ground electronic states of Al^{+}(C_{5}H_{5}N) and its neutral species were determined to be ^{1}A_{1} and ^{2}B_{1} in C_{2}_{v} symmetry. The aluminum-pyridine stretching frequencies were measured to be 278 cm^{−1}in the ionic ^{1}A_{1} state and 366 cm^{−1}in the neutral ^{2}B_{1} state. It is unusual that the metal-ligand stretching frequency of a cation complex is much smaller than that of the corresponding neutral species. This anomaly will be discussed in terms of orbital and electrostatic interactions.

aD. Stöckigt, J. Phys. Chem. A 101, 3800 (1997)